OpenStudy (anonymous):

6/x+4 + 4/x+4=-5 the solution set is

OpenStudy (anonymous):

Since the side of the equation containing the x-values is comprised of 2 fractions with the same denominator, it is fairly easy to remove these fractions; just multiply both sides of the equation by this common denominator: (x+4)(6/x+4 + 4/x+4) = (x+4)(-5) This leaves only the numerators of the original fractions on one side, while moving the x-values to the other side: 6 + 4 = -5x -20 Now solve for x: 10 = -5x -20 30 = -5x x = -6